OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 2x + 4y = 8 slope = -1/2; y-intercept = (0, 2) slope = -2; y-intercept = (0, 4) slope = -2; y-intercept = (0, 8) slope = -1/2; y-intercept = (0, 8)

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

slope of \[ax+by=c\] is \[-\frac{a}{b}\]

OpenStudy (anonymous):

so slope is \[-\frac{2}{4}=-\frac{1}{2}\]

OpenStudy (anonymous):

oh hold on i am wrong

OpenStudy (anonymous):

solve it linerally for y...you get 4y = -2x+8 divide each term by 4...you get y = -1/2x + 2 which is your y = mx + b formula....m is your slope and b is your intercept

OpenStudy (across):

\[2x+4y=8\]\[4y=8-2x\]\[y=2-(1/2)x\]Slope: -1/2 Y-intercept: 2

OpenStudy (anonymous):

it is the first one. to find y intercept put x = 0 solve for y you get \[4y=8\] \[y=2\] so it is indeed \[(0,2)\] i was mistaken

OpenStudy (anonymous):

however it is not necessary to solve for y to find either the slope or the y intercept. to find y - intercept set x = 0 solve to find x - intercept put y = 0 and solve